OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -x + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

OpenStudy (goformit100):

y = -2x - 8

OpenStudy (goformit100):

solve it from straigh line equations with one intersept

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = =1/2x + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x